# FoMo-Bench: a multi-modal, multi-scale and multi-task Forest Monitoring Benchmark for remote sensing foundation models

### v2

@@ 6 Experiments
The models include ResNet-50 and DeiT-S [65] backbones pre-trained on the COCO dataset (as shown in the Table). We also include a Faster R-CNN with a DINOv2 backbone [45] fully finetuned. The FoMo-Net backbone is frozen and the evaluation is performed with a randomly initialized Faster R-CNN head.
We pretrain PolyNet simultaneously on NeonTree and ReforesTree, following which the in-domain evaluation (PolyNet) is performed on each of these datasets. For the out-of-domain evaluation (PolyNet), we pretrain the backbone on one dataset while evaluating on the other.
@@
Semantic segmentation of point clouds: Processing point cloud representations from LiDAR sensors aims at better understanding the geometry of trees at large scale. Classifying each point, i.e. segmenting a point cloud, can be used to distinguish trees, delineate their components (e.g. trunk, branches or leaves) or evaluate their traits (e.g. height, species or biomass). FoMo-Bench includes NeonTree and FORinstance datasets for binary and multi-class point cloud segmentation respectively. For these datasets, we use the baselines PointNet [50], PointNet++[51] and Point Transformer [71]. The performances are presented in Table 6, reported with the mean intersection over union (mIoU) and the F1-score at micro level.
@@ 7 Discussion
7.1 FoMo-Bench and TalloS
We have evaluated performance on FoMo-Bench for a range of models trained on individual datasets as well as providing initial steps towards foundation models for forest-monitoring applications.
@@ 7.1 FoMo-Bench and TalloS
| NeonTree [68] | Airborne | 1 | F. R-CNN | ResNet-50 | 41.08 | 61.49 () | 38.37 () | 13.72 () |
@@
| F. R-CNN | FoMo-Netm | 18.59 | 43.95 () | 0.57 () | 0.10 () | |||
| ReforesTree [53] | Aerial | 6 | F. R-CNN | ResNet-50 | 41.10 | 64.67 () | 6.83 ( | 2.91 () |
@@
| F. R-CNN | FoMo-Netm | 18.60 | 46.40 () | 0.12 () | 0.02 () |
Most of the presented tasks and datasets, especially for object detection and point cloud segmentation, are highly challenging and pose promising areas for future inquiry. We find that standard computer vision models such as ConvNext and Vision Transformers perform relatively poorly on several datasets, e.g. TalloS and Sen12MS. Specifically, the TalloS training set contains more than a thousand classes (see Table 1) making a single-label classification task challenging, and the multi-label classification one even more so.
In datasets such as TalloS, tree species are distributed unequally around the globe (see Tab. 2(b)), with many classes present only in specific locations. One opportunity to improve performance on such tasks could be to incorporate tree species occurrence maps as biological prior knowledge or to learn such information from the data[42].
7.2 FoMo-Net
We evaluate the performance of FoMo-Net using two pretrained backbones, FoMo-Net1 and FoMo-Netm (see Section 5) across a range of semantic segmentation, classification, and object detection tasks For each downstream task we freeze all parameters of the encoder and train a task-specific head, unless specified otherwise. For the classification tasks, we train a linear layer on top of the frozen encoder. Similarly, for the segmentation tasks we extract the encoded tokens using FoMo-Net, upsample to the input resolution and perform the pixel based classification using a multilayer convolutional segmentation head, similar to [5]. The evaluation of object detection tasks also exploits upsampled features of FoMo-Net as inputs of a randomly initialized Faster R-CNN head. In this benchmark, we report the raw performance of our models without employing task-specific strategies such as weighted loss or oversampling techniques.
Segmentation tasks: The reported performances in Table 3 highlight that FoMo-Net pre-training provides relevant semantic information. The FoMo-Net1 offers a robust foundation model baseline competing with most specialized methods for semantic segmentation, even though its finetuning process is not tailored for dense predictions. Even more, its ability to perform well on both aerial and satellite data, despite significant differences in GSD, demonstrates FoMo-Net’s versatility across multiple scales.
| Dataset | Modalities | #C | Method | #P (Mil.) | mIoU | F1 micro |
| NeonTree [68] | Airborne | 1 | PointNet | 62.12 () | 76.63 () | |
| PointNet++ | 1.45 | 55.28 () | 71.20 () | |||
| Point Transf. | 4.59 | 57.70 () | 73.15 () | |||
| FORinstance [49] | Aerial | 7 | PointNet | 35.65 () | 52.56 () | |
| PointNet++ | 1.45 | 33.00 () | 49.57 () | |||
| Point Transf. | 4.19 | 22.97 () | 37.13 () |
Classification tasks: The results presented in Table 4 indicate that the pre-training framework helped to learn significant features for LULC and tree species multi-label classification tasks, though it underperforms supervised approaches on most tasks. Remarkably, the linear evaluation setup of FoMo-Net1, with only 2.000 trainable parameters achieves SoTA performance on TalloS. Even more, when fully finetuned, FoMo-Net1 surpasses the respective ImageNet pretrained ViT architecture (denoted as FoMo-Net1-Finetuned in Table 4). It is worth noting that FoMo-Netm performs consistently worse than FoMo-Net1 in all tasks, in stark contrast to the common practice of learning a dedicated linear transformation per sensor.
Detection tasks: As shown in Table 5, both object detection tasks are challenging. Finetuning using a frozen FoMo-Net backbone performs worse than specialized models but competes with the well-established DINOv2 model with the backbone fully finetuned.
Next Steps: We have seen that our proposed FoMo-Net foundation model for forest monitoring shows promising results on some datasets but leaves considerable room for improvement. Optimizing the pre-training scheme, as well as enhancing its architecture to process non-grid like structures, e.g. LiDAR point clouds, may be especial priorities for future work. This latter direction is achievable by generating appropriate tokens from point clouds and a respective reconstruction projection to the point cloud space. Finally, it may be useful to combine the pre-training approach in FoMo-Net with supervised approaches, for a flexible combination method that can be used under variable computational constraints and with variable amounts of labeled data.
@@ 8 Conclusion
This work presents FoMo-Bench, a multi-modal, multi-scale and multi-task benchmark, geographically distributed worldwide, providing a robust and reliable evaluation framework for future forest monitoring foundation models. To further expand the task and spatial coverage of FoMo-Bench, we introduce TalloS, an extensive globally distributed dataset encompassing multi-temporal and multi-modal samples for tree species classification. Building on these contributions, we craft a foundation model, FoMo-Net, designed for forest-monitoring applications. FoMo-Net leverages a pre-training framework designed to process a flexible range of spectral bands and be applicable for a wide range of tasks in remote sensing. Our experimental evaluation of FoMo-Net on the FoMo-Bench tasks establishes the first baseline for general-purpose forest monitoring, while identifying opportunities and challenges for this promising field.
Appendix 0.A Additional details for FoMo-Net
0.A.1 FoMo-Net pretraining pseudo-code
The pseudocode of the FoMo-Net pre-training pipeline is summarized in Algorithm 1.
Appendix 0.B Evaluation settings of FoMo-Bench
In this section, we outline the training settings employed in FoMo-Bench. By default, we adhere to the training/test splits provided by the original dataset providers unless explicitly stated otherwise. Henceforth, we will denote as supervised models, all models trained end to end in a supervised fashion. All experiments in FoMo-Bench were conducted using a single Nvidia GPU, unless specified otherwise.
0.B.1 FoMo-Net settings
We pretrain both versions of FoMo-Net, namely FoMo-Net1 and FoMo-Netm, for 80 epochs using random resized crop augmentations at 128128 pixels. We set the learning rate to and the weight decay to The training utilized 4 Nvidia A100 GPUs.
0.B.2 Multi-label classification
In this section, we will provide the training settings for our multi-label classification for each dataset. All models have been trained for 20 epochs, using the AdamW optimizer. The experiments including a linear evaluation of FoMo-Net did not use weight decay, unless specified otherwise. We pretrain PolyNet for 50 epochs until the loss reaches a plateau with a learning rate of and a weight decay of . Similar to FoMo-Net we use a cosine learning rate decay schedule.
TalloS: To create the TalloS spatial splits, we divide the world map into 3-degree cells and then randomly allocate them to the train, validation, and test sets to reduce the spatial autocorrelation. Next, to ensure maximum coverage of classes within the training set, we transfer grid cells from the validation and test sets, which contain classes absent in the training set, into the training set. We train all models with a learning rate set to .We set the weight decay to with the exception of the FoMo-Net models, where we did not use weight decay.
RapidAI4EO[2]: We train on RapidAI4EO using a custom nomenclature relevant to forest monitoring. In particular, we aggregate artificial surfaces as one class. We include the following forest-related classes:
1.
Broad-leaved forest
2.
Coniferous forest
3.
Mixed forest
4.
Natural grassland
5.
Moors and heathland
6.
Sclerophyllous vegetation
7.
Transitional woodland/shrub
8.
Sparsely vegetated areas
9.
Burnt areas
10.
Glaciers and perpetual snow
Additionally, we include the following crop type classes:
1.
Non-irrigated arable land
2.
Permanently irrigated land
3.
Rice fields
4.
Vineyards
5.
Fruit trees and berry plantations
6.
Olive groves
7.
Pastures
8.
Annual crops associated with permanent crops
9.
Complex cultivation patterns
10.
Land principally occupied by agriculture, with significant areas of natural vegetation
11.
Agro-forestry areas
Finally, we include one class for water bodies and one for wetlands. RapidAI4EO is spatially organized in directories. We select two areas, i.e. 18E, 27E, as a test set. We used a learning rate of and a weight decay of for all models, with the exception of the FoMo-Net linear evaluation where we used a learning rate of without weight decay.
BigEarthNet-MM[59]: To fully finetune FoMo-Net we train for 50 epochs using a learning rate of and set the weight decay to . For the linear evaluation we set the learning rate to with no weight decay. The training of the rest of the supervised models, including PolyNet, uses a learning rate of and weight decay of .
Sen12MS[57]: For the FoMo-Net linear evaluation, we set the learning rate to . For the supervised training we set the learning rate to and the weight decay to for all models.
0.B.3 Semantic segmentation of grid-based data
All models have been trained for 20 epochs using the AdamW optimizer and a cosine learning rate decay schedule. In the following paragraphs we will describe the training settings for each segmentation dataset. FoMo-Net is adapted to a per-pixel classification scheme by building upon the embeddings of each patch of the input. The extracted features using FoMo-Net have dimensions: , where is the embedding size, is the patch size, and are respectively the height and width dimensions. Subsequently, we upsample these representations to match the input size and feed them into an autoencoder, composed of an encoder with 2 convolutional layers and 2 deconvolutional layers. The final predictions are generated using a convolution layer. No weight decay has been used for FoMo-Net finetuning unless specified otherwise.
ForestNet[29]: We upsample the input to 160160 pixels and train all models using a learning of and a weight decay of for the supervised models (no weight decay for FoMo-Net).
FiveBillionPixels[64]: We separate the scenes that will be used for train and testing and construct 120120 pixel tiles. An upsampled version, i.e. 128128 pixels, is used to train the models with a learning rate of for all models. For the supervised models we use a weight decay of .
Waititu[33]: We designate a single scene defined as the test set. From each scene in the dataset, we generate tiles of size 224224 pixels. We finetune FoMo-Net using a learning rate of with no weight decay, while the fully supervised models with a learning rate of and weight decay of .
Spekboom[17]: Similarly, we hold out 3 scenes for the test set and generate tiles for all sets. We employ a learning rate of for all models and a weight decay of for the supervised models.
Woody[35]: In a similar fashion as with the rest of the UAV datasets i.e. Waititu, and Spekboom, we hold out 3 scenes to serve as our test sets and create tiles of 224224. We train FoMo-Net using a learning rate of , while downsampling the input image to 6464. For the supervised models we use a learning rate of and a weight decay of .
FLAIR#1[20]: We use the original splits provided by the authors. The the FoMo-Net training, we downsample the input to 128128 pixels to speed up training and reduce memory requirements. We set the learning rate to . On the other hand, for the supervised models we keep the original resolution and set the learning rate to and the weight decay to .
FLAIR#2[18]: Given that the labels of the test set were not released at the time of the development, we held out 10% of the aerial scenes from the main dataset to serve as our test set. Similarly to FLAIR#1, we downsample the input size to 6464 pixels for the training of FoMo-Net while using a learning rate of . Since FoMo-Net tokenizes each channel independently, computational complexity is proportional to the number of input channels. For the supervised models, we used a learning rate of and a weight decay of , without downsampling the input image. Except for the UPerNet method where we downsample the input to 128128 pixels and set the learning rate to and the weight decay to .
0.B.4 Object detection
The training settings of all models for the object detection tasks on the NeonTree [68] and ReforesTree [53] datasets are provided in Table 7. The compared methods have either used a stochastic gradient descent (SGD) optimizer or an AdamW optimizer [41]. We also report the intersection over union (IoU) and confidence thresholds, selected within the validation splits, and applied on the predictions over the test splits. During training, we used horizontal flip, vertical flip and Gaussian noise data augmentations. We used the released splits for NeonTree, and we propose our own spatial splits for ReforesTree which are reproducible with the released code (see associated data-loader and script to create tiles from raw data). All experiments have been performed on a single NVIDIA RTX 8000 GPU.
Since DINOv2 [45] and FoMo-Net backbones are transformer-based methods, their last hidden states have been rearranged to squared dimensions and upsampled using bilinear interpolation to the input sizes. These features are processed by an autoencoder with 2 convolutional layers and followed by 2 transposed convolutional layers. A final convolutional layer generates a 32 channels tensor to fit the Faster R-CNN heads and perform the object detection tasks.
In the presented experiments, both the FoMo-Net and PolyNet backbones are frozen during the finetuning of the Faster R-CNN heads, whereas the DINOv2 backbone is fully finetuned with its corresponding heads.
| Dataset | Modality | #C | Method | Backbone | #P (Mil.) | #Epochs | Batch size | Optimizer | LR | Decay | Weight decay | IoU thresh. | Conf. thresh. |
| NeonTree [68] | Airborne | 1 | F. R-CNN | ResNet-50 | 41.08 | 200 | 64 | SGD | Exponential | 0.3 | 0.3 | ||
| RetinaNet | ResNet-50 | 36.13 | 200 | 32 | AdamW | Exponential | 0.5 | 0.3 | |||||
| YOLOS | DeiT-S | 30.65 | 300 | 10 | AdamW | Cosine | - | 0.1 | 0.99 | ||||
| F. R-CNN | PolyNet | 14.50 | 200 | 64 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | DINOv2 | 94.36 | 400 | 16 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | FoMo-Net | 18.59 | 200 | 10 | SGD | Cosine | - | 0.1 | 0.0 | ||||
| ReforesTree [53] | Aerial | 6 | F. R-CNN | ResNet-50 | 41.10 | 200 | 64 | SGD | Exponential | 0.3 | 0.3 | ||
| RetinaNet | ResNet-50 | 36.21 | 300 | 32 | AdamW | Exponential | 0.3 | 0.2 | |||||
| YOLOS | DeiT-S | 30.65 | 300 | 10 | AdamW | Exponential | 0.1 | 0.6 | |||||
| F. R-CNN | PolyNet | 14.52 | 200 | 64 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | DINOv2 | 94.38 | 500 | 16 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | FoMo-Net | 18.60 | 400 | 2 | SGD | Cosine | - | 0.1 | 0.0 |
| Dataset | Modalities | #C | Method | #P (Mil.) | #Epochs | Batch size | LR |
| NeonTree [68] | Airborne | 1 | PointNet | 300 | 16 | ||
| PointNet++ | 1.45 | 500 | 16 | ||||
| Point Transf. | 4.59 | 500 | 16 | ||||
| FORinstance [49] | Aerial | 7 | PointNet | 300 | 16 | ||
| PointNet++ | 1.45 | 400 | 8 | ||||
| Point Transf. | 4.49 | 400 | 2 |
0.B.5 Semantic segmentation of point clouds
The training settings of all models for the semantic segmentation tasks on point clouds applied to the NeonTree [68] and FORinstance [49] datasets are provided in Table 8. The compared methods have all used the AdamW optimizer [41], with an exponential decay of and a random sampling of 1024 points per sample. During training, we used random Jitter and random rotations on the x, y, z axis as data augmentation. Note that the PointNet++ method only used random Jitter and random rotations on the z axis. We used the released splits for both NeonTree and FORinstance datasets. Both datasets provide point clouds with one million points or more per file; thus they have been spatially split in voxels containing 100,000 points at most (see associated script to create sub-point clouds from raw data).
Appendix 0.C TalloS additional information
0.C.1 TalloS - ERA5
We pair TalloS samples with the following variables from ERA5:
1.
volumetric soil water, layers 1-4
2.
skin reservoir content
3.
skin temperature
4.
soil temperature, levels 1-4
5.
surface pressure
6.
total precipitation
7.
2m temperature
8.
2m dewpoint temperature
9.
surface net solar radiation
10.
surface net thermal radiation
11.
surface solar radiation downwards
12.
surface thermal radiation downwards

### v3

@@ 6 Experiments
The models include ResNet-50 and DeiT-S (Touvron et al. 2021) backbones pre-trained on the COCO dataset. We also include a Faster R-CNN with a DINOv2 backbone (Oquab et al. 2024) fully finetuned. FoMo-Net is augmented with a randomly initialized Faster R-CNN head for finetuning. We pretrain PolyNet simultaneously on NeonTree and ReforesTree. In-domain evaluation (PolyNet) is then conducted on each dataset separately. For the out-of-domain evaluation (PolyNet), we pretrain the backbone on one dataset while evaluating on the other.
@@
Semantic segmentation of point clouds: Processing point cloud representations from LiDAR sensors aims at better understanding the geometry of trees at large scale. FoMo-Bench includes NeonTree and FORinstance datasets for binary and multi-class point cloud segmentation respectively. For these datasets, we use the baselines PointNet (Qi et al. 2017a), PointNet++ (Qi et al. 2017b) and Point Transformer (Zhao et al. 2021). These experiments are presented in Appendix E.
@@ 7 Discussion
| NeonTree (Weinstein et al. 2021) | Airborne | 1 | F. R-CNN | ResNet-50 | 41.08 | 61.49 () | 38.37 () | 13.72 () |
@@
| F. R-CNN | FoMo-Netm | 88.55 | 45.12 () | 1.1 () | 0.24 () | |||
| ReforesTree (Reiersen et al. 2022) | Aerial | 6 | F. R-CNN | ResNet-50 | 41.10 | 64.67 () | 6.83 ( | 2.91 () |
@@
| F. R-CNN | FoMo-Netm | 88.57 | 46.38 () | 0.20 () | 0.04 () |
FoMo-Bench: Interestingly, traditional CNNs consistently rank among the top-performing models in both classification and segmentation tasks (Tab. 3, 2). ResNet50 exceptional performance on TalloS is striking, with FoMo-Net1 coming close. TalloS high number of classes along with their skewed distribution make it an extreme multilabel classification problem, reflecting the complexities of global-scale forest monitoring. Tree species are distributed unequally around the globe (see Fig. 2(b)), with many classes present only in specific locations. Incorporating tree species occurrence maps as biological prior knowledge could prove helpful. PolyNet’s performance raises interesting research questions on the utilization of the semantic information of forest monitoring datasets for the creation of foundation models. With the exception of TalloS, where it is severely outperformed by FoMo-Net, PolyNet shows good in-domain predictive skill, including attaining the best performance for Sen12MS. FoMo-Net pretraining yielded promising results, even though it is still outperformed by specialized models. Notably, in some cases it surpasses the original ViT, particularly in TalloS and RapidAI4EO. As discussed in Sec. 5, digesting such diverse information simultaneously might benefit from a larger encoder. Despite the lack of built-in functionality in ViT for processing hierarchical features in dense prediction, FoMo-Net’s performance in segmentation tasks (Table 3) is highly competitive to supervised specialized architectures e.g. UPerNet. Although the supervised baselines in FoMo-Bench set a high bar for comparison, they are not directly comparable to models like FoMo-Net. While supervised-learning models can excel when trained on large datasets, they struggle when applied to different tasks and are fundamentally limited to the specific tasks they were trained on. FoMo-Net pushes the boundaries of this emerging field, establishing a basis for the development of future foundation models. Another key finding is the superiority of FoMo-Net1 over FoMo-Netm. Contrary to intuition, a single shared projection to the embedding space outperforms dedicated projections for each spectral band. This enhances FoMo-Net’s flexibility as new modalities can rely on the shared embedding. We compare FoMo-Net with other foundation models in Appendix F.
Limitations and next steps: While FoMo-Bench offers strong baselines across a broad range of forest monitoring tasks, assessing the true capability of flexible foundation models requires more than just performance metrics. Given the diversity in input modalities we need methods to evaluate how well these flexible models recognize the unique characteristics of each modality, their inter-modal relationships, and the features of the Earth’s surface across multiple scales. Also, as FoMo-Net tokenizes spectral bands independently the amount of tokens per sample increases proportionally to the number of spectral bands, creating a computational bottleneck due to the memory requirements and the quadratic complexity of attention posing challenges for scalability and making optimization a priority. Given PolyNet’s potential, combining FoMo-Net with supervised pre-training might result in a flexible pre-training method usable under variable computational constraints and with variable amounts of labeled data. Finally, extending FoMo-Net to non-grid data, e.g. point clouds, could prove valuable.
@@ 8 Conclusion
This work presents FoMo-Bench, a global benchmark, providing a robust evaluation framework for future forest monitoring flexible foundation models. To further expand the tasks and spatial coverage of FoMo-Bench, we introduce TalloS, a global, multi-temporal and multi-modal dataset for tree species classification. Building on these contributions, we craft a pre-training framework to construct flexible foundation models i.e. FoMo-Net, and provide an initial assessment on FoMo-Bench identifying current limitations and potential opportunities for this promising field.
Acknowledgments
This work has received funding from the IVADO program on “AI, Biodiversity and Climate Change”, the Canada CIFAR AI Chairs program and the European Union’s Horizon Europe research and innovation program under grant agreement No. 101130544 (ThinkingEarth).